In this equation, is X positive, negative, or zero? x+(-8)=-5
Answer by ikleyn(52778)

.
In this equation, is X positive, negative, or zero? x+(-8)=-5
~~~~~~~~~~~~~~~~~

x + (-8) = -5.

Add 8 to both side. You will get

x = -5 + 8,

x = 3.

Answer. x = 3. x is positive three.